Given the function f(x) = 2xsquared - x-1
(f) List the y - intercepts, if there is any, of the graph of f .
(g) Find all point(s) at which f'(x) =0?
(h) Determine the intervals on which the graph of f (x) is increasing and the intervals on which
it is decreasing.
(i) Use the First Derivative Test to identify the maximum and/or minimum point(s) of f (x) .
(j) Use the Second Derivative Test to determine the nature of concavity of f (x) .
1
Expert's answer
2020-06-04T20:33:12-0400

f) Solution

At y-intercept "x=0"

Therefore "y=2(0)^2-(0)-1"

"y=-1"

Answer: The y -intercept is "(0,-1)"


g) Solution

"f(x)=2x^2-x-1"

"f'(x)=4x-1"

At critical point "f'(x)=0"

"4x-1=0"

"4x=1"

"x=0.25"

Answer: At critical point "x=0.25"


h) Solution

From the critical point, we have intervals "(-\u221e,0.25)" and "(0.25,\u221e)"

Let's take "-1" in interval "(-\u221e,0.25)"

"f'(-1)=4(-1)-1"

"-4-1=-5"

Therefore "(-\u221e,0.25)" is a decreasing interval since "f'(x)" is negative.

Let's take "1" in interval "(0.25,\u221e)" ;

"f'(1)=4(1)-1"

"=3"

Therefore "(0.25,\u221e)" is an increasing interval since "f'(x)" is positive.

Answer: "f(x)" is decreasing on interval "(-\u221e,0.25)" and increasing on interval "(0.25,\u221e)"


i) Solution

At maximum or minimum "f'(x)=0"

"f'(x)=4x-1"

"4x-1=0"

"x=0.25"

Substitute "-1,0.25" and "1" in "f(x)" to determine whether "x=0.25" is maximum or minimum

"f(-1)=2(-1)^2-(-1)-1"

"f(-1)=2"

"f(0.25)=2(0.25)^2-(0.25)-1"

"f(0.25)=-1.125"

"f(1)=2(1)^2-(1)-1"

"f(1)=0"

Answer: From above it is clear that "f(x)" is minimum at "x=0.25"


j) Solution

"f'(x)=4x-1"

"f''(x)=4"

Answer: Since the second derivative is positive the function is concave up.